# PsyDI: Towards a Personalized and Progressively In-depth Chatbot for Psychological Measurements

### v1

@@ Abstract
In the field of psychology, traditional psychology testing methods are often criticized for their static nature, lack of customization, and diminished engagement, while in-depth counseling-style assessments remain inaccessible to the general public. The challenge of quantifying psychological traits further complicates these assessments. Despite advancements in psychology measurements using large language models (LLMs), most approaches still rely on single-round Q&A interactions. Addressing this gap, we introduce PsyDI, a personalized and progressively in-depth chatbot designed for psychological measurements, exemplified by its application in the Myers-Briggs Type Indicator (MBTI) framework. PsyDI leverages user-related multi-modal information and engages in customized, multi-turn interactions to more accurately discern the user’s MBTI type. To overcome the challenge of quantifying absolute values of psychological traits, PsyDI applies a novel training paradigm that implicitly estimates unmeasurable variables by learning the rank of proxy variables, culminating in a robust score model for MBTI assessments. This model enables PsyDI to conduct comprehensive and precise MBTI measurements through multi-turn interactions within a unified estimation context based on LLMs. Through various experiments, we validate the efficacy of both the score model and the PsyDI pipeline, demonstrating its potential to serve as a general framework for psychological assessment. Additionally, the online version, PsyDI Online, has revealed a significant user engagement with over 2,000 visits, collecting extensive data tagged with MBTI types and multi-modal information, which facilitates further insightful analysis.
@@ 1 Introduction
Recent progress in general-purpose foundation models, such as large language models (LLMs)(Achiam et al., 2023b; Anthropic, 2023; Touvron et al., 2023) and vision language models (VLMs)(Li et al., 2023; Team et al., 2023; Liu et al., 2024) has shown that artificial intelligence (AI) systems have capabilities to chat, reason, and incorporate relevant context to realize naturalistic interactions. Their advancements provide an opportunity to expand the forms of psychological assessments with AI and make it customized for each user. Traditional techniques often utilize self-report scales and questionnaires for surveys and analysis (Cohen et al., 1996). While this approach can quickly cover a large population, it suffers from relatively weak data reliability and is unable to conduct targeted tests for specific individuals. On the other hand, professional psychological interviews and counseling are not easily accessible to the general public. Correspondingly, due to building upon billions of human natural language knowledge, an LLM-powered system can provide user-specific interaction experience in the psychological dialogue. These agents have the potential to uncover the uncertainties on the surface and dig out the underlying information of different users, then summarizing and analyzing these information based on some professional psychological knowledge injected in prompts or fine-tuning data.
@@
Several recent works attempt to combine neural network techniques with clinical knowledge to improve the performance and generalization on mental disease counseling (Toleubay et al., 2023) and emotional support (Buechel et al., 2018; Cheng et al., 2023) tasks. To enhance the richness of knowledge and the logical reasoning, various methods (Zhang et al., 2024a) have introduced LLMs, fine-tuning them with a variety of specialized psychological data to make them more suitable for these scenarios. Other approaches (Tu et al., 2024a) have adopted role-playing dialogues, guiding LLMs to provide a more immersive psychological dialogue experience. And some of these methods offer an online chat entry for trial, significantly lowering the barrier to entry and promoting wider adoption among the population.
@@
Figure 1: Comparison of traditional psychological testing and the PsyDI framework. (left) Traditional testing, which relies on general questions, can only discern external behavioral traits of the user. This often leads to erroneous MBTI assessments due to controversial external behaviors, such as the INFP person exhibiting some ENFP behaviors. (right) In contrast, the PsyDI framework poses questions based on familiar life scenarios, such as topics like friendship or workplace, and gradually delves into the user’s internal cognitive functions through their external behaviors. This comprehensive approach allows for a more accurate identification of the underlying causes of controversial behaviors and provides a precise MBTI assessment.
@@
Despite these advancements, the existing approaches remain confined to the realm of single-turn textual question-answering (Nori et al., 2023). These methods struggle to maintain style and logic consistency across multi-turn dialogues and fail to offer accurate and stable quantitation measurements over multiple instances for the same user (as shown in Table 1). In this paper, we undertake a rigorous investigation to devise a comprehensive framework that not only offers an engaging interactive experience but also delivers reliable quantifiable analysis outcomes (Figure 1). Our exploration starts from a formalization of psychological assessments within the context of multi-turn decision-making processes (Section 2.2). Specifically, psychological assessments with LLMs are conceived as a multi-turn interaction process designed to evaluate specific discrete types or continuous scores of the test taker. These assessments should be presented in a manner that is both engaging and succinct, ensuring its acceptability to the test-taker. We encapsulate these considerations into the following principles: 1 A superior psychological assessment should actively and ethically gather and construct the user portrait from multi-modal question-answering formats. 2 It should represent a complex skill whose optimal result is highly dependent on the user context, including cultural, linguistic, and individual factors to minimize the biased impact.. 3 It should exhibit comparable or better quantifiable results than existing scales and questionnaires, with rigorous validation and standardization processes.
@@
Based on the above discussed principle, we introduce a comprehensive psychological assessment framework named PsyDI, and validate its effectiveness on one of the most popular personality assessment: Myers-Briggs Type Indicator (MBTI) (Myers et al., 1962). This framework are primarily composed of two innovative designs: a pipeline and a score model. Firstly, recognizing the scarcity of complete multi-turn MBTI assessment dialogue data, we deem it impractical to design an end-to-end dialogue system. Consequently, drawing inspiration from widely-used psychological scales, we adopt a multi-stage dialogue paradigm and design an MBTI profile to select topics for each stage and quantitatively monitor the user’s MBTI predilection. This design leverages the powerful in-context learning abilities of LLMs to construct an engaging multi-turn chat within each stage, while updating the MBTI profile at the end of each stage, effectively decoupling these two components. Furthermore, to ensure accurate adaptive update of MBTI profile, we propose a series of automated data generation scheme and neural network architectures to train a score model, akin to the concept of preference-based reinforcement learning (Wirth et al., 2017). All these components are integrated into a multi-modal and interactive chatbot.
@@
To demonstrate the effectiveness of our proposed framework, we initially conduct several quantitative experiments and ablations about several key designs, including comparisons between various LLMs and our score models across diverse datasets. Subsequently, we release an online accessible version of PsyDI and collect data from over 3,000 participants. Based on de-identified data, we perform a series of analysis and visualizations, examining both specific questions and the whole test process. These examinations validate the framework’s proficiency in dissecting the underlying characteristics of users. Concurrently, it provides qualitative evidence of PsyDI’s capacity for continuous refinement in the assessment of the user’s MBTI. Additionally, we extend PsyDI to the emotion analysis scenario to validate its transferability (Appendix E.4).
The main contributions of this paper can be summarized as follows:
@@
  • We introduce PsyDI, an innovative AI agent framework designed to transcend the limitations of static psychological assessments (e.g. MBTI) through a multi-modal, user-oriented dynamic interaction.
@@
  • To enhance the validity and explanability of PsyDI, we propose a simple yet effective data generation and score model training schemes, with a novel training method that implicitly estimates unmeasurable variables by learning the ranking of proxy variables.
@@
  • Experimental findings derived from quantitative metrics and qualitative real-world user test, encompassing 3,000 participants, substantiate the potential of PsyDI as a general psychological interaction agent.
@@ 2 Problem Formulation
2.1 The Background of MBTI
The unique combination of these four dimensions results in sixteen distinct personality types, each with its own cognitive patterns and behavioral tendencies. This heterogeneity has contributed to the growing popularity of the MBTI in recent years, with over 1.207 billion test administrations recorded by 16Personalities  [^1] https://www.16personalities.com/. However, it is imperative to acknowledge that the MBTI is based on Carl Jung’s theory of psychological types (Jung, 1923). Jung’s theory postulated a quaternary of cognitive functions—thinking, feeling, sensation, and intuition—each with an orientation of either extroversion or introversion, resulting in eight dominant functions, Te (Extroverted Thinking), Ti (Introverted Thinking), Fe (Extroverted Feeling), Fi (Introverted Feeling), Se (Extroverted Sensing), Si (Introverted Sensing), Ne (Extroverted Intuition), and Ni (Introverted Intuition). The constitution of an individual’s personality is delineated by the hierarchical preference for these dominant functions. According to the theory of Jung, it is improbable for a person to exhibit high predilections for both extraversion and introversion within the same cognitive function, as human cognitive resources are finite. This principle leads to the emergence of sixteen common cognitive function sequences, aligning congruently with the sixteen MBTI types. For example, the cognitive function sequence of INFP is: Fi, Ne, Si, Te, Fe, Ni, Se, Ti. Consequently, the adjudication of MBTI typologies inherently entails the scrutiny of one’s cognitive processes.
2.2 Problem Formulation
@@
To better model the MBTI evaluation, we conceptualize the interaction between LLM agents and the test taker (user) as a Markov Decision Process (MDP) (Sutton & Barto, 2018). This formulation allows us to systematically optimize the sequence of inquiries posed to the user, with the aim of enhancing the accuracy, efficiency and engagement of MBTI evaluations. Formally, the MDP is articulated as =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\}$. The states $s$ in the observable state space $\mathcal{S}$ are defined as $s=\{p_{1},\ldots,p_{n}\}$, where $p_{i}$ is the $i$-th textual statement of the user (e.g. I like to talk to different people and share my stories). Additionally, the state space includes a special terminated state $s_{\text{exit}}$ representing the user’s decision to exit the evaluation process.
@@
Throughout the process, the agent needs to determine to either persistently delve deeper into user’s cognitive preferences through multi-turn iterative question-answer (QA), or to terminate the process and render a forecasted MBTI type of the user. While gathering more question-answers may narrow down the user’s authentic MBTI, it is imperative to acknowledge that an extended engagement may engender user dissatisfaction. Therefore, the action space $\mathcal{A}$ is defined as the union of these two kinds of action set $\mathcal{A}=\mathcal{A}_{q}\cup\mathcal{A}_{t}$. Here each $a$ in $\mathcal{A}_{q}$ represents a textual question $a=\{w_{1},\ldots,w_{L}\}$ and $\mathcal{A}_{t}$ represents the set of termination actions, i.e., all possible MBTI types, where $L$ is the length of the question. Once the user selects an answer, the question together with the answer can be regarded as a more detailed user statement $p^{\prime}$. Consequently, the state transition for the state $s$ is $s^{\prime}=s\oplus p^{\prime}$, where $\oplus$ signifies the concatenation of the current state with the new statement $p^{\prime}$. In the event that the user opts to exit, the state transmutes to $s_{\text{exit}}$.
@@
$r(s_{t},a_{t})=\left\{\begin{aligned} -c&,&a\in\mathcal{A}_{t}&,s^{\prime}=s_{\text{exit}}\\ -|a,m^{*}|-\lambda\cdot t&,&a\in\mathcal{A}_{q}&\end{aligned}\right.$ (1)
@@
3 PsyDI Pipeline
In this section, we present the motivation and detailed design aspects of the PsyDI pipeline. Recognizing the intricate nature of human cognitive functions, the task of collecting quantifiable MBTI labels for machine learning applications presents a significant challenge. Nevertheless, we can draw upon the methodologies employed in classical psychological scales. These scales typically consist of questions meticulously crafted by experts. And they often include specific quantitative metrics to assess the test-taker’s during testing. In line with this approach, we have devised an MBTI profile that dynamically summarizes and records the MBTI predilection derived from the test-taker’s statements. This profile is continuously updated with new personal descriptions or feedback obtained from the test-taker, thereby refining the MBTI profile in real-time.
Leveraging this design, the remainder of the PsyDI is structured into two distinct phases: statement selection and multi-turn chat. The former is dedicated to identifying statements with the greatest potential for in-depth questioning, as elaborated in Section 3.2. Then, the multi-turn chat phase is designed to engage users in a series of iterative questions based on the selected statement, fostering an environment that encourages them to provide statements that accurately reflect their cognitive processes while also mitigating the risk of user attrition (Section 3.3). Finally, PsyDI will ascertain the user’s MBTI upon achieving a confident estimation.
@@ 3.1 MBTI Profile
Similar to the quantitative metrics used in psychological scales, we initially establish a kind of estimation of the user’s MBTI type, derived by their statements. In practice, we adopt the psychology profile commonly used in psychological assessments Watson & Clark (1994). This component serves to document the user’s MBTI predilections and subsequently directs other elements of PsyDI in determining the most efficacious strategy for conducting multi-turn dialogues. Specifically, we design a MBTI Profile that documents the individual scores for each trait under assessment. It can translate the user’s statement into quantifiable indicators that accurately reflect their MBTI predilections, while ensuring uniformity and cross-user comparability. Although the MBTI is composed of four distinct dimensions, it is recognized that these dimensions are not entirely independent. The impact of a single statement on each MBTI type is complex, influenced by both the dimensions and the underlying cognitive functions. Therefore, we assign a predilection to each type in MBTI profile, thereby encapsulating the possibility that the current user aligns with that particular MBTI type. As the profile is iteratively updated with fresh statements, the estimated predilections of all the MBTI types are dynamically recalibrated to reflect the evolving understanding of the user.
In the interactive process of PsyDI, the system can only observe the statements provided. Therefore, the likelihood that a user is classified under each MBTI type is determined by how well each statement indicates that the user belongs to a specific MBTI type, which we define as Representativeness. To achieve this, we train a Score Model $\mathcal{F}_{mbti}(\cdot)$ to estimate the representativeness of a statement under each mbti for each dimension, the result of which should be a 4-dimension vector. Then the value of the statement under a MBTI type is defined as the proportion of its representativeness within the distribution of all the statements for the MBTI type:
$\mathcal{V}(p)=\frac{|\{p^{\prime}\in P|\sum_{p\in s}\mathcal{F}(p^{\prime})\leq\sum_{p\in s}\mathcal{F}(p)\}|}{|P|}$ (2)
where $P$ is the comprehensive set of all statements made by users who have been classified with the corresponding MBTI type. $\mathcal{F}(p)$ is the representativeness of statement $p$ while $\mathcal{V}(p)$ represents the value. This value can be calculated under each MBTI separately. The entire training process of Score Model will be introduced in Section 4.
@@
Initialization of MBTI Profile: In the initialization phase, we compute the average representativeness of all known statements attributed to the user. This operation is performed for each MBTI type, allowing us to construct the user’s initial profile:
$\mathcal{V}_{init}(s)=\frac{|\{p^{\prime}\in P|\sum\mathcal{F}(p^{\prime})\leq\sum\bar{\mathcal{F}}(s)\}|}{|P|}$ (3)
Update of MBTI Profile: In the update phase, the incorporation of new statements serves as a means to refine the accuracy of the MBTI profile. Therefore, after obtaining the MBTI representativeness for a new statement, we proceed to recalibrate the MBTI profile according to both the new statement’s representativeness and the established MBTI profile:
$\mathcal{V}_{\text{update}}(s\cup\{p\})=\mathcal{V}(s)+f(\mathcal{V}(s))\cdot\mathcal{V}(p)$ (4)
@@
where $p$ is the latest statement, while $f(\cdot)$ is the non-linear mapping function designed to judiciously regulate the growth weights in accordance based on the current profile.

### v3

@@ Abstract
In the field of psychology, traditional assessment methods, such as standardized scales, are frequently critiqued for their static nature, lack of personalization, and reduced participant engagement, while counseling evaluations are often inaccessible to the general public. The complexity of quantifying psychological traits further limits these methods. Despite advances with large language models (LLMs), many still depend on single-round Question-and-Answer interactions. To bridge this gap, we introduce PsyDI, a personalized and progressively in-depth chatbot designed for psychological measurements, exemplified by its application in the Myers-Briggs Type Indicator (MBTI) framework. PsyDI leverages user-related multi-modal information and engages in customized, multi-turn interactions to provide personalized, easily accessible measurements, while ensuring precise MBTI type determination. To address the challenge of unquantifiable psychological traits, we introduce a novel training paradigm that involves learning the ranking of proxy variables associated with these traits, culminating in a robust score model for MBTI measurements. The score model enables PsyDI to conduct comprehensive and precise measurements through multi-turn interactions within a unified estimation context. Through various experiments, we validate the efficacy of both the score model and the PsyDI pipeline, demonstrating its potential to serve as a general framework for psychological measurements. Furthermore, the online deployment of PsyDI has garnered substantial user engagement, with over 3,000 visits, resulting in the collection of numerous multi-turn dialogues annotated with MBTI types, which facilitates further research. The source code for the training and web service components is publicly available as a part of OpenDILab at: https://github.com/opendilab/PsyDI
@@ 1 Introduction
Recent progress in general-purpose foundation models, such as large language models (LLMs)(Anthropic, 2023; Touvron et al., 2023) and vision language models (VLMs)(Li et al., 2023; Team et al., 2023; Liu et al., 2024) has shown that artificial intelligence (AI) systems have capabilities to chat, reason, and incorporate relevant context to realize naturalistic interactions. Their advancements provide an opportunity to expand the forms of psychological measurements with AI and make it customized for each user. Traditional assessments often utilize self-report standardized scales and questionnaires for surveys and analysis (Cohen et al., 1996). While this approach can quickly cover a large population, it suffers from relatively weak data reliability and is unable to conduct targeted tests for specific individuals. On the other hand, professional psychological interviews and counseling are not easily accessible to the general public. Correspondingly, due to building upon billions of human language knowledge, an LLM-powered system can provide user-specific interaction experiences in the psychological dialogue. These agents have the potential to uncover the uncertainties on the surface and dig out the underlying information of different users, then summarizing and analyzing these information based on some professional psychological knowledge injected in prompts or fine-tuning data.
@@
Several recent works attempt to combine neural network techniques with clinical knowledge to improve the performance and generalization on mental disease counseling (Toleubay et al., 2023) and emotional support (Buechel et al., 2018; Cheng et al., 2023) tasks. To enhance the richness of knowledge and the logical reasoning, various methods (Tu et al., 2024) have introduced LLMs, fine-tuning them with a variety of specialized psychological data to make them more suitable for these scenarios. Other approaches (Yang et al., 2024) have adopted role-playing dialogues, guiding LLMs to provide a more immersive psychological dialogue experience. And some of these methods offer an online chat entry for trial, significantly lowering the barrier to entry and promoting wider adoption among the population.
@@
Figure 1: Comparison of traditional psychological assessment and the PsyDI framework. (left) Traditional assessment, which relies on general questions, can only discern external behavioral traits of the user. This often leads to erroneous MBTI measurements due to controversial external behaviors, such as the INFP person exhibiting some ENFP behaviors. (right) In contrast, the PsyDI framework poses questions based on familiar life scenarios, such as topics like friendship or workplace, and gradually delves into the user’s internal cognitive functions through their external behaviors. This comprehensive approach allows for a more accurate identification of the underlying causes of controversial behaviors and provides a precise MBTI measurement.
@@
Despite these advancements, the existing works remain confined to the realm of single-turn textual question-and-answer (Nori et al., 2023) (Q&A). These methods struggle to maintain style and logic consistency across multi-turn dialogues and fail to offer accurate and stable quantitation measurements over multiple instances for the same user. In this paper, we undertake a rigorous investigation to devise a comprehensive framework that not only offers an engaging interactive experience but also delivers reliable quantifiable analysis outcomes (Figure 1). Our exploration starts from a formalization of psychological measurements within the context of multi-turn decision-making processes (Section 2). Specifically, psychological measurements with LLMs are conceived as a multi-turn interaction process designed to evaluate specific discrete types or continuous scores of the test-taker. These measurements should be presented in a manner that is both engaging and succinct, ensuring its acceptability to the test-taker. We encapsulate these considerations into the following principles: 1 A superior psychological measurement should, akin to the meticulous approach of human experts in psychological typology, actively and ethically gather and construct the user portrait from multi-modal question-answering formats. 2 It should represent a complex skill whose optimal result is highly dependent on the user context, including cultural, linguistic, and individual factors to minimize the biased impact. (Altarriba & Santiago-Rivera, 1994) 3 It should provide solutions for psychological traits that are difficult to quantify, while achieving results that are equivalent to or superior than those yielded by current scale-based assessments, with rigorous validation and standardization processes.
@@
Based on the above discussed principle, we introduce a comprehensive psychological measurement framework named PsyDI, and demonstrate its effectiveness on one of the most popular personality measurement: Myers-Briggs Type Indicator (MBTI) (Myers et al., 1962). This framework are composed of two innovative designs: progressively in-depth pipeline and score model. Firstly, recognizing the scarcity of complete multi-turn MBTI measurement dialogue data, we deem it impractical to design an end-to-end dialogue system. Thus, we adopt a progressively multi-phase dialogue paradigm. Drawing inspiration from widely-used psychological scales, we also design an MBTI profile to select topics from user’s statements for each phase and quantitatively monitor the user’s MBTI predilection. This design leverages the powerful in-context learning abilities of LLMs to construct an engaging multi-turn chat within each phase, while selecting the dialogue topic and updating the MBTI profile at the start and the end of each phase, effectively decoupling these two components. Furthermore, to refine the updates of the profile, we propose a score model training technique, which ranks MBTI indicativeness by leveraging the accuracy of how well statements are predicted to align with a specific MBTI type as a proxy variable. Additionally, this technique employs a ranking loss function to enhance robustness against noise. We then detail a series of automated data generation scheme and neural network architectures to train a score model, akin to the concept of preference-based reinforcement learning (Wirth et al., 2017). All these components are integrated into a multi-modal and interactive chatbot.
@@
To demonstrate the effectiveness of this framework, we initially conduct several quantitative experiments and ablations about several key designs, including comparisons between various LLMs and our score models across diverse datasets (Section 5.2 and Section 5.3). Subsequently, we release an online accessible version of PsyDI and collect data from over 3,000 participants. Based on de-identified data, we perform a series of analysis and visualizations, examining both specific questions and the whole test process. These examinations validate the framework’s proficiency in dissecting the underlying characteristics of users. Concurrently, it provides qualitative evidence of PsyDI’s capacity for continuous refinement in the measurement of the user’s MBTI. Additionally, we extend PsyDI to the emotion analysis scenario to validate its transferability (Appendix E.4). The main contributions of this paper can be summarized as follows:
@@
  • We introduce PsyDI, a novel AI agent framework designed to transcend the limitations of static psychological measurements (e.g. MBTI) through a multi-modal, personalized, progressively in-depth pipeline.
@@
  • To enhance the validity and explanability of PsyDI, we propose a simple yet effective data generation schemes for training the score model, with a novel optimization method that implicitly estimates unmeasurable variables by learning the ranking of proxy variables.
@@
  • Experimental findings derived from quantitative metrics and qualitative real-world user tests, encompassing 3,000 participants, substantiate the potential of PsyDI as a general psychological interaction agent.
@@ 2 Problem Formulation
We initiate our analysis by modelling the measurement of psychological indicators based on the characteristics of MBTI, as detailed in the Appendix A. We conceptualize the interaction between LLM agents and the test taker (user) as a Markov Decision Process (MDP) (Sutton & Barto, 2018). This formulation allows us to systematically optimize the sequence of inquiries posed to the user, with the aim of enhancing the accuracy, efficiency and engagement of MBTI evaluations. Formally, the MDP is articulated as =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\}$. The states $s$ in the observable state space $\mathcal{S}$ are defined as $s=\{p_{1},\ldots,p_{n}\}$, where $p_{i}$ is the $i$-th textual statement of the user (e.g. I like to talk to different people and share my stories). Additionally, the state space includes a special terminated state $s_{\text{exit}}$ representing the user’s decision to exit the evaluation process.
@@
Throughout the process, the agent needs to determine to either persistently delve deeper into user’s cognitive preferences through multi-turn iterative Q&A, or to terminate the process and render a forecasted MBTI type of the user. We define MBTI as $m$, a four-dimensional vector composed of elements 1 and -1, which symbolizes the orientation (E/I, N/S, F/T, J/P) of the MBTI along its four dimensions. For instance, the MBTI type INFP is represented by the vector: $m=[-1,1,1,-1]$. While gathering more question-answers may narrow down the user’s authentic MBTI, it is imperative to acknowledge that an extended engagement may engender user impatient. Therefore, the action space $\mathcal{A}$ is defined as the union of these two kinds of action set $\mathcal{A}=\mathcal{A}_{q}\cup\mathcal{A}_{t}$. Here each $a$ in $\mathcal{A}_{q}$ represents a textual question $a=\{w_{1},\ldots,w_{L}\}$ and $\mathcal{A}_{t}$ represents the set of termination actions, i.e., all possible MBTI types, where $L$ is the length of the question. Once the user selects an answer, the question together with the answer can be regarded as a more detailed user statement $p^{\prime}$. Consequently, the state transition for the state $s$ is $s^{\prime}=s\oplus p^{\prime}$, where $\oplus$ signifies the concatenation of the current state with the new statement $p^{\prime}$. In the event that the user opts to exit, the state transmutes to $s_{\text{exit}}$.
@@
$r(s_{t},a_{t})=\left\{\begin{aligned} -c&,&a\in\mathcal{A}_{q}&,s^{\prime}=s_{\text{exit}}\\ -|a,m^{*}|-\lambda\cdot t&,&a\in\mathcal{A}_{t}&\end{aligned}\right.$ (1)
@@
3 PsyDI Progressively In-depth Pipeline
In this section, we present the motivation and detailed design of the PsyDI pipeline. The task of collecting quantifiable MBTI labels for machine learning applications presents a significant challenge. Nevertheless, we can draw upon the methodologies employed in classical psychological scales, which often include specific quantitative metrics to assess the test-taker’s during testing. In line with this approach, we have devised an MBTI profile that dynamically summarizes and records the MBTI predilection derived from the test-taker’s statements, as described in Section 3.1. This profile is continuously updated with new self-reported statements or feedback obtained from the summary of the multi-turn chat, thereby refining the MBTI profile in real-time. Leveraging this design, the remainder of PsyDI is structured into two distinct phases: statement selection and multi-turn chat. The former is dedicated to identifying statements with the greatest potential for questioning (Section 3.2). Then, the multi-turn chat phase is designed to engage users in iterative questions based on the selected statement, fostering an environment that encourages them to provide statements that accurately reflect their cognitive processes while also reducing the chance that they might stop participating (Section 3.3). Finally, PsyDI will ascertain the user’s MBTI upon achieving a confident estimation. The entire workflow of PsyDI is illustrated in Figure 2. For more details on the pipeline, please refer to Appendix D.
@@ 3.1 MBTI Profile
Similar to the quantitative metrics used in psychological scales, we initially establish a kind of estimation of the user’s MBTI type, derived by their statements. In practice, we adopt the psychology profile commonly used in psychological measurements (Watson & Clark, 1994). This component serves to document the user’s MBTI predilections and subsequently directs other modules of PsyDI in determining the most efficacious strategy for conducting multi-turn dialogues. Specifically, we design an MBTI Profile that documents the individual scores for each trait under measurement. It can translate the user’s statement into quantifiable indicators that accurately reflect their MBTI predilections, while ensuring uniformity and cross-user comparability.
Considering the non-independence among the four dimensions of the MBTI, we assign a predilection to each type, encapsulating the likelihood that the user aligns with that specific MBTI type. The advantage of this approach is that it allows us to independently model the impact of a statement on each type. Furthermore, as the profile is continually updated with new statements, the estimated predilections for all MBTI types are dynamically adjusted to accurately reflect the evolving understanding of the user’s personality traits.
The state influences the likelihood of a user being classified under each MBTI type. This classification is based on the degree to which each statement reflects the characteristics of a specific MBTI type, a metric we refer to as Indicativeness. To quantify it, we train a Score Model $\mathcal{F}_{m}(\cdot)$ to estimate the indicativeness of a statement under MBTI $m$. For the definition and the entire training process of Score Model, please refer to Section 4. Then the value of a statement within an MBTI type is determined by its quantile point, reflecting its level of indicativeness within the overall distribution of statements associated with MBTI type $m$:
$\mathcal{V}_{m}(p)=\frac{|\{p^{\prime}\in P_{m}|\mathcal{F}_{m}(p^{\prime})\leq\mathcal{F}_{m}(p)\}|}{|P_{m}|}$ (2)
where $P$ is the entire set of all statements made by users with MBTI $m$. $\mathcal{F}_{m}(p)$ is the indicativeness of statement $p$ while $\mathcal{V}_{m}(p)$ represents the value. For example, a statement with a value of 0.75 represents that 75% of the statements with the same MBTI in the dataset has an indicativeness less than this statement.
@@
Initialization of MBTI Profile: For this, we compute the average indicativeness of all known statements attributed to the user. This operation is performed for each MBTI, allowing us to construct the user’s initial profile, where $s$ represents for state and ¯$\bar{\mathcal{F}}_{m}(\cdot)$ is the average of the indicativeness of all the statements in $s$:
$\mathcal{V}^{0}_{m}(s)=\frac{|\{p^{\prime}\in P_{m}|\mathcal{F}_{m}(p^{\prime})\leq\bar{\mathcal{F}}_{m}(s)\}|}{|P_{m}|}$ (3)
Update of MBTI Profile: During the update of MBTI profile, the incorporation of new statements serves as a way to refine the accuracy of the MBTI profile. This update process implicitly accounts for the interdependencies among the MBTI types, as the scoring model, trained to recognize these relationships, updates the scores in response to new statement. The impact of each new statement on the profile is contingent upon the initial estimates; for instance, if a user’s profile strongly favors INFP, additional statements may not significantly alter this inclination. Consequently, the update process involves a linear mapping function that adjusts the MBTI profile based on the indicativeness of the new statement relative to the existing profile:
$\mathcal{V}_{m}(s\cup\{p_{new}\})=\mathcal{V}_{m}(s)+f(\mathcal{V}_{m}(s))\cdot\mathcal{V}_{m}(p_{new})$ (4)
@@
where $p_{new}$ is the latest statement, while $f(\cdot)$ is the non-linear mapping function designed to judiciously regulate the growth weights in accordance based on the current profile.